# Fundamental limitations on the recoverability of quantum processes

### v1

@@ C.5 Proof of additivity of channel entropy
With
$X$ $:=D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\sigma_{R^{\prime}A}\right),$
$Y$ $:=D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A}\right),$ (117)
and then exploiting the monotonicity of relative entropy under partial trace, the above equation leads to the following inequality
$X$ $\leq Y+D\left(\rho_{R^{\prime}A_{1}}\;\middle\|\;\sigma_{R^{\prime}A_{1}}\right)$
$\leq Y+D\left(\rho_{R^{\prime}B_{2}A_{1}}\;\middle\|\;\sigma_{R^{\prime}B_{2}A_{1}}\right)$
$=Y+D\left(\left(\mathrm{id}_{R}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\otimes\mathrm{id}_{A_{1}}\right)\Psi_{R^{\prime}A_{2}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R}\otimes\mathcal{R}_{A_{2}\rightarrow B_{2}}\otimes\mathrm{id}_{A_{1}}\right)\Psi_{R^{\prime}A_{2}A_{1}}\right)$
$=Y+D\left(\left(\mathrm{id}_{R}\otimes\mathrm{id}_{A_{1}}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\right)\Psi_{R^{\prime}A_{1}A_{2}}\;\middle\|\;\left(\mathrm{id}_{R}\otimes\mathrm{id}_{A_{1}}\otimes\mathcal{R}_{A_{2}\rightarrow B_{2}}\right)\Psi_{R^{\prime}A_{1}A_{2}}\right).$ (118)
Taking supremum on both sides of the above inequality with respect to state $\Psi_{RA_{1}A_{2}}$, we then obtain
$\sup_{\Psi_{RA_{1}A_{2}}}X\leq\sup_{\Psi_{RA_{1}A_{2}}}Y+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right],$ (119)
@@
$D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes\mathcal{R}_{A_{2}\rightarrow B_{2}}\right]$ $\leq\sup_{\Psi_{RA_{1}A_{2}}}Y+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right]$
$=D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\right]+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right].$ (120)
By multiplying with the minus sign on both sides of the above inequality and using the definition of the entropy of the quantum channel, we obtained the desired inequality
@@
[…]
@@
[…]
By applying inequalities (121) and (125), we can conclude that the entropy functional of a quantum channel is additive under tensor product of channels i.e,
[…]
@@ C.6 Entropy gain under positive maps
For a given pair of input and output quantum states $\{\sigma,\rho\}$, where the output state $\rho$ is always a pure state but the input state could be either a pure or a mixed state, there always exists an unital CP map $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$, such that $\rho=\mathcal{F}(\sigma)$.
@@
Consider a set $\mathtt{K}_{C}=\{\sigma,\mathbbm{1}\}\subset\mathcal{L}\left(C\right)$, and let $\mathcal{A}_{\mathtt{K}_{C}}$ denote the sub-algebra generates by $\mathtt{K}_{C}$. Clearly, $\mathcal{A}_{\mathtt{K}_{C}}$ is a commutative unital $C^{*}$ sub-algebra of $\mathcal{L}(C)$. Now, we define a map $\mathcal{F}:\mathcal{A}_{\mathtt{K}_{C}}\rightarrow\mathcal{L}\left(A\right)$ by $\mathcal{F}(\sigma)=\rho$, $\mathcal{F}(\sigma^{n})=\rho$, for all $n\in\mathbb{N}/\{0\}$, and $\mathcal{F}(\mathbbm{1})=\mathbbm{1}_{A}$, and extend it by linearity to the whole space. As $\sigma$ is positive and hence, hermitian, it is clear that the map $\mathcal{F}$ is a ${}^{\ast}-\text{homomorphism}$ which also implies that $\mathcal{F}$ is a positive map. As $\mathcal{A}_{\mathtt{K}_{C}}$ is a commutative unital $C^{*}$ algebra and $\mathcal{L}\left(A\right)$ is an unital $C^{*}$ algebra, we have that $\mathcal{F}$ is completely positive Pau (03). We now employ Arveson’s extension theorem to extend it to $\mathcal{L}\left(C\right)$.
@@
Note that our proof will also hold if we take output state $\rho$ to be a mixed state if it satisfies the following condition: $\forall~{}c_{n}\in\mathbb{C}$ if $\sum_{n}c_{n}\sigma^{n}=0$, then $\sum_{n}c_{n}\rho^{n}$ must vanish. Then, we can define a map $\mathcal{F}:\mathcal{A}_{\mathtt{K}_{C}}\rightarrow\mathcal{L}\left(A\right)$ by $\mathcal{F}(\sigma)=\rho^{n}$, $\mathcal{F}(\sigma^{n})=\rho$, for all $n\in\mathbb{N}/\{0\}$, and $\mathcal{F}(\mathbbm{1})=\mathbbm{1}_{A}$, and extend it by linearity to the whole space.
@@
Let $\rho\in\mathcal{D}\left(A\right)$, $\sigma\in\mathcal{D}\left(C\right)$ and $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ be a completely positive unital map that connects $\rho$ and $\sigma$. Then, the von Neumann entropy of $\rho$ and $\sigma$ are related by the following equation:
$S\left(\rho\right)\geq\norm{\mathcal{F}^{*}(\mathbbm{1}_{A})}_{\infty}S\left(\sigma\right)$ (127)
where, *$\mathcal{F}^{*}:\mathcal{L}\left(A\right)\rightarrow\mathcal{L}\left(C\right)$ is the adjoint of $\mathcal{F}$.
@@
Let $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ denote an unital and completely positive map, then we can define a completely positive trace nonincreasing map ~$\widetilde{\mathcal{F}}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ as follows $\widetilde{F}(\sigma):=\frac{\mathcal{F}(\sigma)}{\norm{\mathcal{F}^{*}(\mathbbm{1})}_{\infty}}$. Moreover, for any unital and completely positive map $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$, and for any positive operator $\sigma\in\mathcal{L}\left(\mathcal{H}_{A}\right)$, the following operator inequality holds (see Wol (12)), $\mathcal{F}\left(\sigma\right)\log\mathcal{F}\left(\sigma\right)\leq\mathcal{F}\left(\sigma\log\sigma\right)$, now, by replacing $\mathcal{F}$ in terms of $\widetilde{\mathcal{F}}$ in the right-hand side of this inequality, we obtain $\mathcal{F}\left(\sigma\right)\log\mathcal{F}\left(\sigma\right)\leq\norm{\mathcal{F}^{*}\left(\mathbbm{1}\right)}_{\infty}\widetilde{\mathcal{F}}(\sigma\log\sigma)$. By taking trace on both sides and using the fact that $\widetilde{F}$ is trace nonincreasing, we then have $-S(\left(\mathcal{F}\left(\sigma\right)\right)\leq-\norm{\mathcal{F}^{*}\left(\mathbbm{1}\right)}_{\infty}S\left(\sigma\right)$, which implies that $S\left(\rho\right)\geq\norm{\mathcal{F}^{*}\left(\mathbbm{1}\right)}_{\infty}S\left(\sigma\right)$.
Note that in principle, there may exist more than one unital CP map connecting $\rho$ to $\sigma$  (which we also proved earlier in the beginning of this section). Thus, this lower-bound can be further tightened as $S(\rho)\geq\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}\right)}_{\infty}\}S\left(\sigma\right)$, where the supremum is taken over all possible $\mathcal{F}$. Furthermore, notice that the value of $\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}\right)}_{\infty}\}$ decides whether the entropy gain is positive or negative. Thus, we can make the following remark: for a completely positive unital map $\mathcal{F}$, entropy gain is non-negative if $\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}\right)}_{\infty}\}\geq 1$, and entropy gain is negative if $\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}\right)}_{\infty}\}<1$. Once we assume that map $\mathcal{F}$ is also trace preserving, which implies that $\mathcal{F}^{*}$ is unital, we then have $\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}\right)}_{\infty}\}=1$, thus by retrieving the well-known result AU (77). ∎
@@
Let $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ be a completely positive map. Then, for all $\sigma\in\mathcal{D}\left({C}\right)$ such that $\mathcal{F}\left(\sigma\right)>0$, the change in entropy after and before applying the map is given by
$\Delta S\left(\sigma,\mathcal{F}\right)\geq\max\left\{D\left(\sigma\|\sigma_{\alpha}\right),D\left(\sigma\|\hat{\sigma}_{\alpha}\right)\right\}$ (128)
where, $\Delta S\left(\sigma,\mathcal{F}\right):=S\left(\mathcal{F}\left(\sigma\right)\right)-S\left(\sigma\right)$ is the entropy gain, $\alpha:=\norm{\mathcal{F}^{*}(\mathbbm{1}_{A})}$, $\sigma_{\alpha}:=\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathcal{F}^{*}\circ\mathcal{F}(\sigma)\right)^{\alpha}$, and $\hat{\sigma}_{\alpha}:=\frac{1}{\alpha}\mathcal{F}^{*}\left(\mathcal{F}(\sigma)\right)^{\alpha}$.
@@
We define a completely positive, trace nonincreasing map ~$\widetilde{\mathcal{F}}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ from map $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ as $\widetilde{F}(\sigma):=\frac{1}{\alpha}\mathcal{F}(\sigma)$, then, by using the definition of adjoint the entropy gain can be written as $\Delta S\left(\sigma,\mathcal{F}\right)=\operatorname{tr}\left(\sigma\log\sigma\right)-\operatorname{tr}\left(\alpha\sigma\widetilde{\mathcal{F}}^{*}\left(\log\mathcal{F}\left(\sigma\right)\right)\right)$. As $\mathcal{F}\left(\sigma\right)$ is a full rank operator by assumption, we can use the fact that for any sub-unital map $\widetilde{\mathcal{F}}$, we have, ~$\widetilde{\mathcal{F}}\left(\log(\sigma)\right)\leq\log\left(\widetilde{\mathcal{F}}\left(\sigma\right)\right)$ if $\mathcal{F}\left(\sigma\right)>0$, to arrive at following inequality $\Delta S\left(\sigma,\mathcal{F}\right)\geq\operatorname{tr}\left(\sigma\log\sigma\right)-\operatorname{tr}\left(\sigma\alpha(\log\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}\left(\sigma\right))\right)$, thus by arriving at the entropy inequality, $\Delta S\left(\sigma,\mathcal{F}\right)\geq D\left(\sigma\left|\right|\sigma_{\alpha}\right)$, where $\sigma_{\alpha}:=\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathcal{F}^{*}\circ\mathcal{F}(\sigma)\right)^{\alpha}$. Note that we can also write entropy gain as $\Delta S\left(\sigma,\mathcal{F}\right)=\operatorname{tr}\left(\sigma\log\sigma\right)-\operatorname{tr}\left(\widetilde{\mathcal{F}}\left(\sigma\right)\log\left(\mathcal{F}\left(\sigma\right)\right)^{\alpha}\right)$, and after this, by following the same arguments as above, we obtain another bound on entropy gain as $\Delta S\left(\sigma,\mathcal{F}\right)\geq D\left(\sigma\left|\right|\hat{\sigma}_{\alpha}\right)$, where $\hat{\sigma}_{\alpha}:=\frac{1}{\alpha}\mathcal{F}^{*}\left(\mathcal{F}(\sigma)\right)^{\alpha}$. Since, both the remainder terms, $D\left(\sigma\|\sigma_{\alpha}\right)$ and $D\left(\sigma\|\hat{\sigma}_{\alpha}\right)$ give a valid bound on entropy gain, so, to obtain a better bound we take a maximum of both the remainder terms. Thus, we have
$\Delta S\left(\sigma,\mathcal{F}\right)\geq\max\left\{D\left(\sigma\|\sigma_{\alpha}\right),D\left(\sigma\|\hat{\sigma}_{\alpha}\right)\right\}.$ (129)
If we take $\mathcal{F}^{*}$ to be unital (or equivalently $\mathcal{F}$ to be a quantum channel), then we have $\alpha=1$, and we obtain a well-known result BDW (16): $\Delta S\left(\sigma,\mathcal{F}\right)\geq D\left(\sigma\|\mathcal{F}^{*}\circ\mathcal{F}(\sigma)\right)$. ∎
We now state some known facts which we are going to use to prove our next result.
@@
    Let $A$ and $B$ be two finite-dimensional Hilbert spaces and let $\mathcal{F}:\mathcal{L}(A)\rightarrow\mathcal{L}(B)$ be a positive linear map such that $\mathcal{F}(\mathbbm{1})\leq\mathbbm{1}$. Then for all $\mathsf{X}>0$ we have the following inequality Wol (12):
@@
    Let $\rho\in\mathcal{L}(A)$ be a quantum state and $\sigma\in\mathcal{L}(A)_{+}$ be a positive semidefinite operator. Then the following limit holds Wil (13):
    $D(\rho\|\sigma)=\lim_{\epsilon\rightarrow 0+}D(\rho\|\sigma+\epsilon\mathbbm{1}).$ (131)
@@
    Let $\rho\in\mathcal{L}(A)$ be a quantum state and $\sigma,\sigma^{{}^{\prime}}\in\mathcal{L}(A)_{+}$ be positive semidefinite operators such that $\sigma\leq\sigma^{{}^{\prime}}$. Then we have the following inequality Wil (13) :
    $D(\rho\|\sigma)\geq D(\rho\|\sigma^{{}^{\prime}}).$ (132)
@@
where $\alpha:=\norm{\mathcal{F}^{*}(\mathbbm{1})}_{\infty}$ and $\rho_{\alpha}:=\frac{1}{\alpha^{\alpha}}\left(\mathcal{F}^{*}\circ\mathcal{F}(\rho)\right)^{\alpha}$.

### v2

@@ C.5 Proof of additivity of channel entropy
Define
$X$ $:=D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\sigma_{R^{\prime}A}\right),$ (124)
$Y$ $:=D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A}\right).$ (125)
That is, $X=Y+D\left(\rho_{R^{\prime}}\;\middle\|\;\sigma_{R^{\prime}}\right)$. Now exploiting the monotonicity of the relative entropy under partial trace, we have
$X$ $\leq Y+D\left(\rho_{R^{\prime}B_{2}A_{1}}\;\middle\|\;\sigma_{R^{\prime}B_{2}A_{1}}\right)$
$=Y+D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\otimes\mathrm{id}_{A_{1}}\right)\Psi_{R^{\prime}A_{2}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{2}\rightarrow B_{2}}\otimes\mathrm{id}_{A_{1}}\right)\Psi_{R^{\prime}A_{2}A_{1}}\right)$
$=Y+D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathrm{id}_{A_{1}}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\right)\Psi_{R^{\prime}A_{1}A_{2}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathrm{id}_{A_{1}}\otimes\mathcal{R}_{A_{2}\rightarrow B_{2}}\right)\Psi_{R^{\prime}A_{1}A_{2}}\right).$ (126)
Taking supremum on both sides of the above inequality with respect to state $\Psi_{R^{\prime}A_{1}A_{2}}$, we then obtain
$\sup_{\Psi_{R^{\prime}A_{1}A_{2}}}X\leq\sup_{\Psi_{R^{\prime}A_{1}A_{2}}}Y+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right],$ (127)
@@
$D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes\mathcal{R}_{A_{2}\rightarrow B_{2}}\right]$ $\leq D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\right]+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right].$ (128)
By multiplying with the minus sign on both sides of the above inequality and using the definition of the entropy of the quantum channel, we obtain the desired inequality
@@
$\geq\sup_{\widetilde{\Psi}}D\left((\mathrm{id}_{R_{1}}\otimes{\mathcal{N}}_{A_{1}\rightarrow B_{1}})\Psi_{R_{1}A_{1}}\otimes(\mathrm{id}_{R_{2}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}})\Psi_{R_{2}A_{2}}\;\middle\|\;(\mathrm{id}_{R_{1}}\otimes{\mathcal{R}}_{A_{1}\rightarrow B_{1}})\Psi_{R_{1}A_{1}}\otimes(\mathrm{id}_{R_{2}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}})\Psi_{R_{2}A_{2}}\right)$
@@
$S\left[{\mathcal{N}}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\right]\leq S\left[{\mathcal{N}}_{A_{1}\rightarrow B_{1}}]+S[{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\right].$ (133)
Combining inequalities (129) and (133), we conclude that the entropy functional of a quantum channel is additive under tensor product of channels, i.e,
@@ C.6 Entropy gain under positive maps
For a given pair of input and output quantum states $\{\sigma\in\mathcal{D}\left(C\right),\rho\in\mathcal{D}\left(A\right\}$, where the output state $\rho$ is always a pure state but the input state could be either a pure or a mixed state, there always exists a unital CP map $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$, such that $\rho=\mathcal{F}(\sigma)$.
@@
Consider a set $\mathtt{K}_{C}=\{\sigma,\mathbbm{1}\}\subset\mathcal{L}\left(C\right)$, and let $\mathcal{A}_{\mathtt{K}_{C}}$ denote the sub-algebra generated by $\mathtt{K}_{C}$. Clearly, $\mathcal{A}_{\mathtt{K}_{C}}$ is a commutative unital $C^{*}$ sub-algebra of $\mathcal{L}(C)$. Now, we define a map $\mathcal{F}:\mathcal{A}_{\mathtt{K}_{C}}\rightarrow\mathcal{L}\left(A\right)$ by $\mathcal{F}(\sigma)=\rho$, $\mathcal{F}(\sigma^{n})=\rho$, for all $n\in\mathbb{N}/\{0\}$, and $\mathcal{F}(\mathbbm{1}_{C})=\mathbbm{1}_{A}$, and extend it by linearity to the whole space. As $\sigma$ is positive and hence, hermitian, it is clear that the map $\mathcal{F}$ is a ${}^{\ast}-\text{homomorphism}$ which also implies that $\mathcal{F}$ is a positive map. As $\mathcal{A}_{\mathtt{K}_{C}}$ is a commutative unital $C^{*}$ algebra and $\mathcal{L}\left(A\right)$ is a unital $C^{*}$ algebra, we have that $\mathcal{F}$ is completely positive Pau (03). We now employ Arveson’s extension theorem to extend it to $\mathcal{L}\left(C\right)$.
@@
Note that our proof will also hold if we take output state $\rho$ to be a mixed state if it satisfies the following condition: $\forall~{}c_{n}\in\mathbb{C}$ if $\sum_{n}c_{n}\sigma^{n}=0$, then $\sum_{n}c_{n}\rho^{n}$ must vanish. Then, we can define a map $\mathcal{F}:\mathcal{A}_{\mathtt{K}_{C}}\rightarrow\mathcal{L}\left(A\right)$ by $\mathcal{F}(\sigma^{n})=\rho^{n}$, for all $n\in\mathbb{N}/\{0\}$, and $\mathcal{F}(\mathbbm{1}_{C})=\mathbbm{1}_{A}$, and extend it by linearity to the whole space.
@@
Let $\rho\in\mathcal{D}\left(A\right)$, $\sigma\in\mathcal{D}\left(C\right)$ and $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ be a completely positive unital map that connects $\rho$ and $\sigma$. Then, the von Neumann entropy of $\rho$ and $\sigma$ are related by the following equation.
$S\left(\rho\right)\geq\norm{\mathcal{F}^{*}(\mathbbm{1}_{A})}_{\infty}S\left(\sigma\right),$ (135)
where $\mathcal{F}^{*}:\mathcal{L}\left(A\right)\rightarrow\mathcal{L}\left(C\right)$ is the adjoint of $\mathcal{F}$.
@@
Let $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ denote an unital and completely positive map, then we can define a completely positive trace nonincreasing map ~$\widetilde{\mathcal{F}}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ as follows ~$\widetilde{\mathcal{F}}(\sigma):=\frac{\mathcal{F}(\sigma)}{\norm{\mathcal{F}^{*}(\mathbbm{1}_{A})}_{\infty}}$. Moreover, for any unital and completely positive map $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$, and for any positive operator $\sigma\in\mathcal{L}\left(C\right)$, the following operator inequality holds (see Wol (12)), $\mathcal{F}\left(\sigma\right)\log\mathcal{F}\left(\sigma\right)\leq\mathcal{F}\left(\sigma\log\sigma\right)=\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}\widetilde{\mathcal{F}}(\sigma\log\sigma)$. By taking trace on both sides and using the fact that $\widetilde{F}$ is trace nonincreasing, we then have $-S(\left(\mathcal{F}\left(\sigma\right)\right)\leq-\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}S\left(\sigma\right)$, which implies that $S\left(\rho\right)\geq\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}S\left(\sigma\right)$.
Note that, in principle, there may exist more than one unital CP map connecting $\rho$ to $\sigma$  (which we also proved earlier in the beginning of this section). Thus, this lower-bound can be further tightened as $S(\rho)\geq\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}\}S\left(\sigma\right)$, where the supremum is taken over all possible maps $\mathcal{F}$ connecting states $\rho$ and $\sigma$. Further, notice that the value of $\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}\}$ decides whether the entropy gain is positive or negative. Thus, we can make the following remark: for a completely positive unital map $\mathcal{F}$, entropy gain is non-negative if $\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}\}\geq 1$, and entropy gain is negative if $\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}\}<1$. Once we assume that map $\mathcal{F}$ is also trace preserving, which implies that $\mathcal{F}^{*}$ is unital, we then have $\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}\}=1$, thereby retrieving the well-known result of AU (77). ∎
@@
Let $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ be a completely positive map and let the entropy gain of $\sigma\in\mathcal{D}\left({C}\right)$ under the map $\mathcal{F}$ be defined as $\Delta S\left(\sigma,\mathcal{F}\right):=S\left(\mathcal{F}\left(\sigma\right)\right)-S\left(\sigma\right)$. Then, for all $\sigma\in\mathcal{D}\left({C}\right)$ such that $\mathcal{F}\left(\sigma\right)>0$, we have
$\Delta S\left(\sigma,\mathcal{F}\right)\geq\max\left\{D\left(\sigma\;\middle\|\;\sigma_{\alpha}\right),D\left(\sigma\;\middle\|\;\hat{\sigma}_{\alpha}\right)\right\},$ (136)
where $\alpha:=\norm{\mathcal{F}^{*}(\mathbbm{1}_{A})}_{\infty}$, $\sigma_{\alpha}:=\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathcal{F}^{*}\circ\mathcal{F}(\sigma)\right)^{\alpha}$, and $\hat{\sigma}_{\alpha}:=\frac{1}{\alpha}\mathcal{F}^{*}\left(\mathcal{F}(\sigma)\right)^{\alpha}$.
@@
We first define a completely positive, trace nonincreasing map ~$\widetilde{\mathcal{F}}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ from map $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ as $\widetilde{F}(\sigma):=\frac{1}{\alpha}\mathcal{F}(\sigma)$, where $\alpha:=\norm{\mathcal{F}^{*}(\mathbbm{1}_{A})}_{\infty}$. Now using the definition of adjoint, the entropy gain can be written as $\Delta S\left(\sigma,\mathcal{F}\right)=\operatorname{tr}\left(\sigma\log\sigma\right)-\operatorname{tr}\left(\alpha\sigma\widetilde{\mathcal{F}}^{*}\left(\log\mathcal{F}\left(\sigma\right)\right)\right)$. As $\mathcal{F}\left(\sigma\right)$ is a full rank operator by assumption, we can use the fact that for any sub-unital map $\widetilde{\mathcal{F}}$, we have ~$\widetilde{\mathcal{F}}\left(\log(\sigma)\right)\leq\log\left(\widetilde{\mathcal{F}}\left(\sigma\right)\right)$ if $\mathcal{F}\left(\sigma\right)>0$. Using this, we get $\Delta S\left(\sigma,\mathcal{F}\right)\geq\operatorname{tr}\left(\sigma\log\sigma\right)-\operatorname{tr}\left(\sigma\alpha(\log\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}\left(\sigma\right))\right)$, which implies $\Delta S\left(\sigma,\mathcal{F}\right)\geq D\left(\sigma\;\middle\|\;\sigma_{\alpha}\right)$, where $\sigma_{\alpha}:=\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathcal{F}^{*}\circ\mathcal{F}(\sigma)\right)^{\alpha}$. Note that we can also write entropy gain as $\Delta S\left(\sigma,\mathcal{F}\right)=\operatorname{tr}\left(\sigma\log\sigma\right)-\operatorname{tr}\left(\widetilde{\mathcal{F}}\left(\sigma\right)\log\left(\mathcal{F}\left(\sigma\right)\right)^{\alpha}\right)$. Then, by following the similar arguments as above, we obtain another bound on entropy gain as $\Delta S\left(\sigma,\mathcal{F}\right)\geq D\left(\sigma\;\middle\|\;\hat{\sigma}_{\alpha}\right)$, where $\hat{\sigma}_{\alpha}:=\frac{1}{\alpha}\mathcal{F}^{*}\left(\mathcal{F}(\sigma)\right)^{\alpha}$. Since both the remainder terms, $D\left(\sigma\;\middle\|\;\sigma_{\alpha}\right)$ and $D\left(\sigma\;\middle\|\;\hat{\sigma}_{\alpha}\right)$, are valid lower bounds on the entropy gain, we have
$\Delta S\left(\sigma,\mathcal{F}\right)\geq\max\left\{D\left(\sigma\;\middle\|\;\sigma_{\alpha}\right),D\left(\sigma\;\middle\|\;\hat{\sigma}_{\alpha}\right)\right\}.$ (137)
This completes the proof of the lemma. ∎
If we take $\mathcal{F}^{*}$ to be unital (or equivalently $\mathcal{F}$ to be a quantum channel) in the above lemma, then we have $\alpha=1$, and we obtain a well-known result BDW (16): $\Delta S\left(\sigma,\mathcal{F}\right)\geq D\left(\sigma\|\mathcal{F}^{*}\circ\mathcal{F}(\sigma)\right)$. We now state some known facts which we will be using to prove our next results.
@@
    Let $A$ and $B$ be two finite-dimensional Hilbert spaces and let $\mathcal{F}:\mathcal{L}(A)\rightarrow\mathcal{L}(B)$ be a positive linear map such that $\mathcal{F}(\mathbbm{1}_{A})\leq\mathbbm{1}_{B}$. Then for all $\mathsf{X}>0$ we have the following inequality Wol (12):
@@
    Let $\rho\in\mathcal{D}(A)$ be a quantum state and $\sigma\in\mathcal{L}(A)_{+}$ be a positive semidefinite operator. Then the following limit holds Wil (13):
    $D\left(\rho\;\middle\|\;\sigma\right)=\lim_{\epsilon\rightarrow 0+}D\left(\rho\;\middle\|\;\sigma+\epsilon\mathbbm{1}_{A}\right).$ (139)
@@
    Let $\rho\in\mathcal{D}(A)$ be a quantum state and $\sigma,\sigma^{\prime}\in\mathcal{L}(A)_{+}$ be positive semidefinite operators such that $\sigma\leq\sigma^{\prime}$. Then we have the following inequality Wil (13) :
    $D\left(\rho\;\middle\|\;\sigma\right)\geq D\left(\rho\;\middle\|\;\sigma^{\prime}\right).$ (140)
In the following, we provide a twofold generalization of Lemma 10: (1) we consider positive maps instead of completely positive maps, (2) we drop the requirement that $\mathcal{F}(\sigma)>0$.
@@
where $\alpha:=\norm{\mathcal{F}^{*}(\mathbbm{1}_{B})}_{\infty}$ and $\rho_{\alpha}:=\frac{1}{\alpha^{\alpha}}\left(\mathcal{F}^{*}\circ\mathcal{F}(\rho)\right)^{\alpha}$.